OpenStudy (anonymous):

In 1990, the life expectancy of males in a certain country was 69.9 years. In 1995, it was 72.8 years. Let E represent the number of years since 1990 E(t)= _ t+ _ ???? (round to the nearest tenth) Use the function to predict the life expectancy of males in 2006 E(16)= ??

OpenStudy (anonymous):

I know there were 5 years since 1990 to 1995, and I know that there is 16 years between 1990 and 2006, I just do not know how to solve this or write it the way they want it

OpenStudy (jamesj):

Look at the problem we just solved together and mimic the calculations. It is exactly the same. I also suggest you do not write E(16) for 2006; instead, keep the variable equal to the year, E(2006).

OpenStudy (anonymous):

You know that E(0) is 69.9 years. That will help you get your equation. If \[E(t) = at + b\], plug in t=0: \[E(0) = a(0) + b = 69.9\], so b = 69.9. Next, you know that E(5) = 72.8 [E(5) being the life expectancy in 1995]. Use this to find a: \[E(t) = at + 69.9\]\[E(5) = a(5) + 69.9 = 72.8\], so you can solve for a as well. Then, you just need to let t=16! \[E(16) = a(16) + 69.9\], which will yield your answer.

OpenStudy (anonymous):

I had to step out for a minute. thanks you guys